An electric dipole with two charges +5uc and -5uc are separated by distance of 10cm is placed in uniform electric field of intensity 200v/m . Calculate the torque acting on dipole if excess of dipole makes an angle 30° with speed.


1
Expert's answer
2021-03-16T11:36:29-0400
"\tau=pE\sin{30}\\\tau=qlE\sin{30}\\\tau=(5\cdot10^{-6})(0.1)(200)\sin{30}=5\cdot10^{-5}\ Nm"